The sum of two consecutive integers is 99
Answer by nerdybill(7384)

The sum of two consecutive integers is 99
.
Let x = first integer
then
x+1 = second integer
.
x + x+1 = 99
2x + 1 = 99
2x = 98
x = 49 (first integer)
.
second integer:
x+1 = 49+1 = 50
.
The two integers are 49 and 50